How would you sovle this problem?
Find the measure of an angle whose supplement measures 5 times its measure.

Found 2 solutions by rfer, Sunny Day:
Answer by rfer(16322)

180/6=30 degrees

Answer by Sunny Day(15)

Let the angle be x.
Then its supplement is 5x.
SInce x and 5x are supplementary,
x +5x = 180
6x = 180
x = 180/6 = 30
so the angle is 30 degrees